\begin{thebibliography}{25}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Abdou et~al.(2021)Abdou, Kulmizev, Hershcovich, Frank, Pavlick, and
  S{\o}gaard}]{abdou-etal-2021-language}
Mostafa Abdou, Artur Kulmizev, Daniel Hershcovich, Stella Frank, Ellie Pavlick,
  and Anders S{\o}gaard. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.conll-1.9} {Can language
  models encode perceptual structure without grounding? a case study in color}.
\newblock In \emph{Proceedings of the 25th Conference on Computational Natural
  Language Learning}, pages 109--132, Online. Association for Computational
  Linguistics.

\bibitem[{Bender and Koller(2020)}]{bender-koller-2020-climbing}
Emily~M. Bender and Alexander Koller. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.463} {Climbing
  towards {NLU}: {On} meaning, form, and understanding in the age of data}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association
  for Computational Linguistics}, pages 5185--5198, Online. Association for
  Computational Linguistics.

\bibitem[{Bhakthavatsalam et~al.(2020)Bhakthavatsalam, Anastasiades, and
  Clark}]{DBLP:journals/corr/abs-2005-00660}
Sumithra Bhakthavatsalam, Chloe Anastasiades, and Peter Clark. 2020.
\newblock \href {http://arxiv.org/abs/2005.00660} {Genericskb: {A} knowledge
  base of generic statements}.
\newblock \emph{CoRR}, abs/2005.00660.

\bibitem[{Chella(2015)}]{Chella2015}
Antonio Chella. 2015.
\newblock \href {https://doi.org/10.1007/978-3-319-15021-5_10} {\emph{A
  Cognitive Architecture for Music Perception Exploiting Conceptual Spaces}},
  pages 187--203. Springer International Publishing, Cham.

\bibitem[{Forbes et~al.(2019)Forbes, Holtzman, and
  Choi}]{DBLP:conf/cogsci/ForbesHC19}
Maxwell Forbes, Ari Holtzman, and Yejin Choi. 2019.
\newblock \href {https://mindmodeling.org/cogsci2019/papers/0311/index.html}
  {Do neural language representations learn physical commonsense?}
\newblock In \emph{Proceedings of the 41th Annual Meeting of the Cognitive
  Science Society, CogSci 2019: Creativity + Cognition + Computation, Montreal,
  Canada, July 24-27, 2019}, pages 1753--1759. cognitivesciencesociety.org.

\bibitem[{Gajbhiye et~al.(2022)Gajbhiye, Espinosa-Anke, and
  Schockaert}]{gajbhiye-etal-2022-modelling}
Amit Gajbhiye, Luis Espinosa-Anke, and Steven Schockaert. 2022.
\newblock \href {https://aclanthology.org/2022.coling-1.349} {Modelling
  commonsense properties using pre-trained bi-encoders}.
\newblock In \emph{Proceedings of the 29th International Conference on
  Computational Linguistics}, pages 3971--3983, Gyeongju, Republic of Korea.
  International Committee on Computational Linguistics.

\bibitem[{G{\"{a}}rdenfors(2000)}]{DBLP:books/daglib/0006106}
Peter G{\"{a}}rdenfors. 2000.
\newblock \emph{Conceptual spaces - the geometry of thought}.
\newblock {MIT} Press.

\bibitem[{Gordon and Durme(2013)}]{DBLP:conf/cikm/GordonD13}
Jonathan Gordon and Benjamin~Van Durme. 2013.
\newblock \href {https://doi.org/10.1145/2509558.2509563} {Reporting bias and
  knowledge acquisition}.
\newblock In \emph{Proceedings of the 2013 workshop on Automated knowledge base
  construction, AKBC@CIKM 13, San Francisco, California, USA, October 27-28,
  2013}, pages 25--30. {ACM}.

\bibitem[{Gärdenfors(2004)}]{gardenfors2004conceptual}
Peter Gärdenfors. 2004.
\newblock Conceptual spaces as a framework for knowledge representation.
\newblock \emph{Mind and Matter}, 2(2):9--27.

\bibitem[{Gärdenfors(2014)}]{10.7551/mitpress/9629.001.0001}
Peter Gärdenfors. 2014.
\newblock \href {https://doi.org/10.7551/mitpress/9629.001.0001} {\emph{{The
  Geometry of Meaning: Semantics Based on Conceptual Spaces}}}.
\newblock The MIT Press.

\bibitem[{He et~al.(2021)He, Gao, and Chen}]{DBLP:journals/corr/abs-2111-09543}
Pengcheng He, Jianfeng Gao, and Weizhu Chen. 2021.
\newblock \href {http://arxiv.org/abs/2111.09543} {Debertav3: Improving deberta
  using electra-style pre-training with gradient-disentangled embedding
  sharing}.
\newblock \emph{CoRR}, abs/2111.09543.

\bibitem[{Ji et~al.(2019)Ji, Wang, Shi, Zhang, Wang, and
  Yan}]{DBLP:journals/dint/JiWSZWY19}
Lei Ji, Yujing Wang, Botian Shi, Dawei Zhang, Zhongyuan Wang, and Jun Yan.
  2019.
\newblock \href {https://doi.org/10.1162/dint\_a\_00013} {Microsoft concept
  graph: Mining semantic concepts for short text understanding}.
\newblock \emph{Data Intell.}, 1(3):238--270.

\bibitem[{Li et~al.(2023)Li, Xu, Dong, Zheng, Liu, Kong, and
  Sun}]{DBLP:journals/corr/abs-2305-14057}
Lei Li, Jingjing Xu, Qingxiu Dong, Ce~Zheng, Qi~Liu, Lingpeng Kong, and Xu~Sun.
  2023.
\newblock \href {https://doi.org/10.48550/arXiv.2305.14057} {Can language
  models understand physical concepts?}
\newblock \emph{CoRR}, abs/2305.14057.

\bibitem[{Liu et~al.(2022{\natexlab{a}})Liu, Eisenschlos, Cole, and
  Collier}]{liu-etal-2022-ever}
Fangyu Liu, Julian Eisenschlos, Jeremy Cole, and Nigel Collier.
  2022{\natexlab{a}}.
\newblock \href {https://aclanthology.org/2022.aacl-short.27} {Do ever larger
  octopi still amplify reporting biases? evidence from judgments of typical
  colour}.
\newblock In \emph{Proceedings of the 2nd Conference of the Asia-Pacific
  Chapter of the Association for Computational Linguistics and the 12th
  International Joint Conference on Natural Language Processing (Volume 2:
  Short Papers)}, pages 210--220, Online only. Association for Computational
  Linguistics.

\bibitem[{Liu et~al.(2022{\natexlab{b}})Liu, Yin, Feng, and
  Zhao}]{liu-etal-2022-things}
Xiao Liu, Da~Yin, Yansong Feng, and Dongyan Zhao. 2022{\natexlab{b}}.
\newblock \href {https://doi.org/10.18653/v1/2022.acl-long.168} {Things not
  written in text: Exploring spatial commonsense from visual signals}.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association
  for Computational Linguistics (Volume 1: Long Papers)}, pages 2365--2376,
  Dublin, Ireland. Association for Computational Linguistics.

\bibitem[{Martin et~al.(2014)Martin, Visalli, Lange, Schlich, and
  Issanchou}]{martin2014creation}
Christophe Martin, Michel Visalli, Christine Lange, Pascal Schlich, and Sylvie
  Issanchou. 2014.
\newblock Creation of a food taste database using an in-home “taste”
  profile method.
\newblock \emph{Food Quality and Preference}, 36:70--80.

\bibitem[{McRae et~al.(2005)McRae, Cree, Seidenberg, and
  McNorgan}]{mcrae2005semantic}
Ken McRae, George~S Cree, Mark~S Seidenberg, and Chris McNorgan. 2005.
\newblock Semantic feature production norms for a large set of living and
  nonliving things.
\newblock \emph{Behavior research methods}, 37(4):547--559.

\bibitem[{Misra et~al.(2022)Misra, Rayz, and
  Ettinger}]{DBLP:journals/corr/abs-2205-06910}
Kanishka Misra, Julia~Taylor Rayz, and Allyson Ettinger. 2022.
\newblock \href {https://doi.org/10.48550/arXiv.2205.06910} {A property
  induction framework for neural language models}.
\newblock \emph{CoRR}, abs/2205.06910.

\bibitem[{Paik et~al.(2021)Paik, Aroca-Ouellette, Roncone, and
  Kann}]{paik-etal-2021-world}
Cory Paik, St{\'e}phane Aroca-Ouellette, Alessandro Roncone, and Katharina
  Kann. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.emnlp-main.63} {{T}he {W}orld
  of an {O}ctopus: {H}ow {R}eporting {B}ias {I}nfluences a {L}anguage
  {M}odel{'}s {P}erception of {C}olor}.
\newblock In \emph{Proceedings of the 2021 Conference on Empirical Methods in
  Natural Language Processing}, pages 823--835, Online and Punta Cana,
  Dominican Republic. Association for Computational Linguistics.

\bibitem[{Paradis(2015)}]{Paradis2015}
Carita Paradis. 2015.
\newblock \href {https://doi.org/10.1007/978-3-319-15021-5_3} {\emph{Conceptual
  Spaces at Work in Sensory Cognition: Domains, Dimensions and Distances}},
  pages 33--55. Springer International Publishing, Cham.

\bibitem[{Patel and Pavlick(2022)}]{patel2022mapping}
Roma Patel and Ellie Pavlick. 2022.
\newblock Mapping language models to grounded conceptual spaces.
\newblock In \emph{International Conference on Learning Representations}.

\bibitem[{S{\o}gaard(2023)}]{DBLP:journals/mima/Sogaard23}
Anders S{\o}gaard. 2023.
\newblock \href {https://doi.org/10.1007/s11023-023-09622-4} {Grounding the
  vector space of an octopus: Word meaning from raw text}.
\newblock \emph{Minds Mach.}, 33(1):33--54.

\bibitem[{Standley et~al.(2017)Standley, Sener, Chen, and
  Savarese}]{DBLP:conf/corl/StandleySCS17}
Trevor Standley, Ozan Sener, Dawn Chen, and Silvio Savarese. 2017.
\newblock \href {http://proceedings.mlr.press/v78/standley17a.html}
  {image2mass: Estimating the mass of an object from its image}.
\newblock In \emph{1st Annual Conference on Robot Learning, CoRL 2017, Mountain
  View, California, USA, November 13-15, 2017, Proceedings}, volume~78 of
  \emph{Proceedings of Machine Learning Research}, pages 324--333. {PMLR}.

\bibitem[{Zhang et~al.(2022)Zhang, Roller, Goyal, Artetxe, Chen, Chen, Dewan,
  Diab, Li, Lin, Mihaylov, Ott, Shleifer, Shuster, Simig, Koura, Sridhar, Wang,
  and Zettlemoyer}]{DBLP:journals/corr/abs-2205-01068}
Susan Zhang, Stephen Roller, Naman Goyal, Mikel Artetxe, Moya Chen, Shuohui
  Chen, Christopher Dewan, Mona~T. Diab, Xian Li, Xi~Victoria Lin, Todor
  Mihaylov, Myle Ott, Sam Shleifer, Kurt Shuster, Daniel Simig, Punit~Singh
  Koura, Anjali Sridhar, Tianlu Wang, and Luke Zettlemoyer. 2022.
\newblock \href {https://doi.org/10.48550/arXiv.2205.01068} {{OPT:} open
  pre-trained transformer language models}.
\newblock \emph{CoRR}, abs/2205.01068.

\bibitem[{Zwarts(2015)}]{Zwarts2015}
Joost Zwarts. 2015.
\newblock \href {https://doi.org/10.1007/978-3-319-15021-5_4} {\emph{Conceptual
  Spaces, Features, and Word Meanings: The Case of Dutch Shirts}}, pages
  57--78. Springer International Publishing, Cham.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{EMNLP2023}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{booktabs}
\usepackage{multirow}

\usepackage{xcolor,caption,subcaption,graphicx}

\title{Cabbage Sweeter than Cake? Analysing the Potential of Large Language Models for Learning Conceptual Spaces}

\author{Usashi Chatterjee, Amit Gajbhiye, Steven Schockaert \\
  CardiffNLP, nCardiff University, UK\\
  \texttt{\{chatterjeeu,gajbhiyea,schockaerts1\}@cardiff.ac.uk} \\}

\begin{document}
\maketitle
\begin{abstract}
The theory of Conceptual Spaces is an influential cognitive-linguistic framework for representing the meaning of concepts. Conceptual spaces are constructed from a set of quality dimensions, which essentially correspond to primitive perceptual features (e.g.\ hue or size). These quality dimensions are usually learned from human judgements, which means that applications of conceptual spaces tend to be limited to narrow domains (e.g.\ modelling colour or taste). Encouraged by recent findings about the ability of Large Language Models (LLMs) to learn perceptually grounded representations, we explore the potential of such models for learning conceptual spaces. Our experiments show that LLMs can indeed be used for learning meaningful representations to some extent. However, we also find that fine-tuned models of the BERT family are able to match or even outperform the largest GPT-3 model, despite being 2 to 3 orders of magnitude smaller.\footnote{Our datasets and evaluation scripts are available at {\url{https://github.com/ExperimentsLLM/EMNLP2023_PotentialOfLLM_LearningConceptualSpace.git}}.}
\end{abstract}

\section{Introduction}
Conceptual spaces \cite{DBLP:books/daglib/0006106} represent concepts in terms of cognitively meaningful features, called quality dimensions. For example, a conceptual space of colour is composed of three quality dimensions, representing hue, saturation and intensity. Conceptual spaces provide an elegant framework for explaining various cognitive and linguistic phenomena \cite{10.7551/mitpress/9629.001.0001}. Within Artificial Intelligence (AI), the role of conceptual spaces is essentially to act as an intermediate representation layer, in between neural and symbolic representations \cite{gardenfors2004conceptual}. As such, conceptual spaces could play a central role in the development of explainable AI systems. Unfortunately, such representations are difficult to learn from data. Most applications of conceptual spaces are thus limited to narrow domains, where meaningful representations can be learned from ratings provided by human participants \cite{Paradis2015,Zwarts2015,Chella2015}. 

In this paper, we explore whether Large Language Models (LLMs) could be used for learning conceptual spaces. This research question is closely related to the ongoing debate about the extent to which Language Models (LMs) can learn perceptually grounded representations \cite{bender-koller-2020-climbing,abdou-etal-2021-language,patel2022mapping,DBLP:journals/mima/Sogaard23}. Recent work seems to suggest this might indeed be possible, at least for the colour domain. For instance, \citet{abdou-etal-2021-language} found that LMs are able to learn representations of colour terms which are isomorphic to perceptual colour spaces. When it comes to predicting the typical colour of objectes, \citet{paik-etal-2021-world} found that the predictions of LMs are heavily skewed by surface co-occurrence statistics, which are unreliable for colours due to reporting bias \cite{DBLP:conf/cikm/GordonD13}, i.e.\ the fact that obvious colours are rarely mentioned in text. However, \citet{liu-etal-2022-ever} found the effects of reporting bias to largely disappear in recent LLMs. These findings suggest that it may now be possible to distill meaningful conceptual space representations from LLMs, as long as sufficiently large models are used. 
However, existing analyses are limited in two ways:
\begin{itemize}
\item Several works have explored the colour domain, and visual domains more generally \cite{DBLP:journals/corr/abs-2305-14057}, but little is known about the abilities of LLMs in other perceptual domains. 
\item Most analyses focus on classifying concepts, e.g.\ predicting colour terms or the materials from which objects are made, rather than on evaluating the underlying quality dimensions. 
\end{itemize}
We address the first limitation by including an evaluation in the taste domain. To address the second limitation, rather than considering discrete labels (e.g. \emph{sweet}), we use LLMs to rank concepts according to the degree to which they have a particular feature (e.g.\ \emph{sweetness}). 

\section{Datasets}
The primary focus of our experiments is on the taste domain, which has not yet been considered in this context, despite having a number of important advantages. For instance, the relevant quality dimensions are well-established and have a linear structure (unlike hue in the colour domain). This domain also seems particularly challenging, as the typical terms which are used to describe taste only apply to extreme cases. For instance, we can assert that grapefruit is bitter and bananas are sweet, but it is less clear how a language model would learn whether chicken is sweeter than cheese. As ground truth, we rely on the ratings that were collected by \citet{martin2014creation}. They rated a total of 590 food items along six dimensions: sweet, salty, sour, bitter, umami and fat. The ratings were obtained by a panel of twelve assessors who were experienced in sensory profiling. They scored the food items during an eight month measurement phase, after having received 55 hours of training in a laboratory. We manually rephrased some of the names of the items in this dataset, to make them more natural. For instance, \emph{cherry (fresh fruit)} was changed to \emph{cherry} and \emph{hake (grilled with lemon juice)} was changed to \emph{grilled hake with lemon juice}.


We complement our analysis in the taste domain with experiments on three basic physical domains: mass, size and height. These were found to be particularly challenging by \citet{DBLP:journals/corr/abs-2305-14057}, with LLMs often failing to outperform random guessing. As the ground truth for mass, we use the household dataset from \citet{DBLP:conf/corl/StandleySCS17}, which specifies the mass of 56 household objects. The original dataset includes images of each object. We removed 7 items which were not meaningful without the image, namely \emph{big elephant}, \emph{small elephant}, \emph{Ivan's phone}, \emph{Ollie the monkey}, \emph{Marshy the elephant}, \emph{boy doll} and \emph{Dali Clock}, resulting in a dataset of 49 objects. 
We treat this problem as a ranking problem. \citet{DBLP:journals/corr/abs-2305-14057} also created a binary classification version of this dataset, which involves judging pairwise comparisons (e.g.\ is a red lego brick heavier than a hammer?).
For size and height, we use the datasets created by \citet{liu-etal-2022-things}. These size and height datasets each consist of 500 pairwise judgements (e.g.\ an ant is larger than a bird). Note that unlike for the other datasets, no complete ranking is provided.

\section{Methods}
We experiment with a number of different models.

\paragraph{Ranking with GPT-3} 
We use GPT-3 models of four different sizes\footnote{The exact model sizes have not been made public, but were estimated to be 350M parameters for \emph{ada}, 1.3B parameters for \emph{babbage}, 6.7B parameters for \emph{curie} and 175B parameters for \emph{davinci}: \url{https://blog.eleuther.ai/gpt3-model-sizes/}.}: \emph{ada}, \emph{babbage}, \emph{curie} and \emph{davinci}. 
To rank items according to a given dimension, we use a prompt that contains the name of that dimension as the final word, e.g.\ for sweetness we could use ``\emph{It is known that [food item] tastes sweet''}. We then use the probability of this final word, conditioned on the rest of the prompt, to rank the item: the higher the probability of \emph{sweet}, the more we assume the item to be sweet. 

\paragraph{Pairwise Comparisons with GPT-3}
To predict pairwise judgements, we consider two approaches. 
First, we again use conditional probabilities. For instance, to predict whether \emph{an ant is larger than a bird}, we would get the conditional probability of \emph{large} in the sentences \emph{an ant is large} and \emph{a bird is large}. If the conditional probability we get from the first sentence is lower than the probability from the second sentence, we would predict that the claim that \emph{an ant is larger than a bird} is false.
Second, we use a prompt that asserts the statement to be true (e.g.\ ``An ant is larger than a bird'') and a prompt that asserts the opposite (e.g.\ ``A bird is larger than an ant''). We compute the perplexity of both statements and predict the version with the lowest perplexity to be the correct one.

\paragraph{Ranking with ChatGPT and GPT-4}
ChatGPT and GPT-4 are more difficult to use than GPT-3 because the OpenAI API does not allow us to compute conditional probabilities for these models. Instead, to use these conversational models, we directly ask them to rank a set of items, using a prompt such as: \emph{Rank the following items according to their size, from the largest to the smallest}, followed by a list of items to be ranked.

\paragraph{Baseline: DeBERTa}
We consider two baselines. First, we use a DeBERTa-v3-large model \cite{DBLP:journals/corr/abs-2111-09543}, which we fine-tuned to predict the commonsense properties of concepts. To this end, we used the extended McRae dataset \cite{mcrae2005semantic} introduced by \citet{DBLP:conf/cogsci/ForbesHC19} and the augmented version of CSLB\footnote{\url{https://cslb.psychol.cam.ac.uk/propnorms}} introduced by \citet{DBLP:journals/corr/abs-2205-06910}. Together, these two datasets contain 19,410 positive and 31,901 negative examples of (concept,property) pairs. We fine-tune the model on these examples using the following prompt: \emph{can <concept> be described as <property>? <MASK>}. The probability that the concept has the property is then predicted using a linear classifier that takes the final-layer embedding of the <MASK> token as input. We use the resulting model for the different evaluations, without any further fine-tuning. 

\paragraph{Baseline: Bi-encoder}
As the second baseline, we use two variants of the bi-encoder model from \citet{gajbhiye-etal-2022-modelling}. First, we use the original BERT-large model from \citet{gajbhiye-etal-2022-modelling} that was trained on data from Microsoft Concept Graph \cite{DBLP:journals/dint/JiWSZWY19} and GenericsKB \cite{DBLP:journals/corr/abs-2005-00660}. However, as these training sets are not specifically focused on commonsense knowledge, we used ChatGPT to construct a dataset of 109K (concept,property) pairs, since no existing dataset of sufficient size and quantity was available. The key to obtain high-quality examples was to ask the model to suggest properties that are shared by several concepts, and to vary the examples that were provided as part of a few-shot prompting strategy. More details on how we collected this dataset using ChatGPT are provided in Appendix \ref{secChatGPTPairs}. We then trained the BERT-large bi-encoder on this dataset.


\section{Experiments}

\begin{table}[t]
\footnotesize
\centering
\begin{tabular}{l @{\hspace{6pt}} l @{\hspace{4pt}} c@{\hspace{7pt}}c@{\hspace{7pt}}c@{\hspace{7pt}}c@{\hspace{7pt}}c@{\hspace{7pt}}c@{\hspace{7pt}} c}
\toprule
&&  \rotatebox{90}{\textbf{Sweet}} &  \rotatebox{90}{\textbf{Salty}} &  \rotatebox{90}{\textbf{Sour}} &  \rotatebox{90}{\textbf{Bitter}} &  \rotatebox{90}{\textbf{Umami}} &  \rotatebox{90}{\textbf{Fatty}}\\
\midrule
\multirow{4}{*}{\rotatebox{90}{\scriptsize\textsc{prompt 1}}}&Ada & 17.5 & 8.5 & 12.2 & 16.4 & 22.5 & 10.7  \\
&Babbage & 19.5 & 51.1 & 20.2 & 22.0 & 22.6 & 16.0  \\
&Curie & 36.0 & 46.3 & 32.8 & 23.2 & 22.6 & 31.7  \\
&Davinci & 55.0 & 63.2 & 33.3 & 27.2 & \textbf{57.0} & 52.0  \\
\midrule
\multirow{4}{*}{\rotatebox{90}{\scriptsize\textsc{prompt 2}}} &Ada & 23.1 & 11.9 & 8.5 & 16.8 & -6.4 & 9.9  \\
&Babbage &27.9 & 55.7 & 19.3 & 23.9 & 29.7 & 34.0  \\
&Curie & 35.4& 47.9 & 30.3 & 22.7 & 25.4 & 37.5  \\
&Davinci & 50.2 & 54.4 & 34.6 & \textbf{28.3} & 49.8 & 42.1  \\
\midrule
&DeBERTa & \textbf{69.1} & \textbf{67.0} & \textbf{43.9} & 24.7 & 34.4 & \textbf{64.0}  \\
&Bi-enc\textsubscript{MSCG+GKB}   & 27.4 & -4.4 & 15.4 & 14.4 & -11.8 & 12.6  \\
&Bi-enc\textsubscript{ChatGPT}  & 60.3 & 47.1 & 40.4 & 9.0 & 40.7 & 40.2  \\
\bottomrule
\end{tabular}
\caption{Ranking using conditional probability (Spearman $\rho$\%). Prompt1: ``[food item] tastes [property]''. Prompt 2: ``it is known that [food item] tastes [property]''.\label{tabTasteRank}}
\end{table}

\begin{table}[t]
\footnotesize
\centering
\begin{tabular}{lcc}
\toprule
\textbf{Item} & \textbf{Gold} & \textbf{Davinci}\\
\midrule
Cracker with Nutella spread & 5 & 324 \\
Chocolate with nut & 15 & 201\\
Sweet pancake with maple syrup & 34 & 279 \\
Fruit cake & 41 & 265 \\
Sweet cookies with chocolate & 60 & 233\\
Cracker with jam & 67 & 268\\
\midrule
Cooked bell pepper & 252 & 8\\
Redcurrant & 282 & 36\\
Radish & 431 & 97\\
Mascarpone cheese & 477 & 38\\
Cooked green cabbage & 512 & 49 \\
Saint-agur Cheese & 584 & 61 \\
\bottomrule
\end{tabular}
\caption{Qualitative analysis of the predictions by \emph{davinci} for sweetness, using prompt 1. The table shows the rank positions of several food items, when ranking the items from the sweetest (rank 1) to the least sweet (rank 590), according to the ground truth and the predictions obtained with the Davinci model. \label{tabQualitativeSweetness}}
\end{table}


\begin{table}[t]
\footnotesize
\centering
\begin{tabular}{l @{\hspace{10pt}} l @{\hspace{5pt}} cccc }
\toprule
&& \textbf{Mass} & \textbf{Mass} & \textbf{Height} & \textbf{Size}\\
&& $\rho$ & \textit{Acc} & \textit{Acc} & \textit{Acc}\\
\midrule
\multirow{4}{*}{\rotatebox{90}{\scriptsize\textsc{cond.\ prob.}}} & Ada & 23.0 & 47.8 & 68.7 & 59.1\\
&Babbage & 48.9 & 80.9 & 67.9 & 76.4\\
&Curie & 30.6 & 65.1 & 77.6 & 86.4\\
&Davinci & 36.2 & 76.8 & 76.4 & 80.4\\
\midrule
\multirow{4}{*}{\rotatebox{90}{\scriptsize\textsc{perplexity}}} & Ada & - & 49.0 & 49.7 & 36.5\\
&Babbage & - & 55.0 & 59.1 & 66.7\\
&Curie & - & 56.6 & 43.3 & 45.5\\
&Davinci & - & 70.8 & 54.7 & 51.1\\
\midrule
\multirow{2}{*}{\rotatebox{90}{\scriptsize\textsc{rank}}} & ChatGPT$^\dagger$ & 28.6 & 68.3 & 89.9 & 84.3\\
&GPT-4$^\dagger$ & \textbf{58.6} & \textbf{84.9} & \textbf{99.1} & \textbf{99.1}\\
\midrule
&DeBERTa & -8.9 & 42.8 & 86.6 & 93.9\\
&Bi-enc\textsubscript{MSCG+GKB}   & 31.1 & 69.2 & 69.7 & 71.9\\
&Bi-enc\textsubscript{ChatGPT}  & 11.8 & 67.6 & 77.2 & 60.3\\
\bottomrule
\end{tabular}
\caption{Results for physical properties, viewed as a ranking problem (mass) and as a pairwise judgment problem (mass, height and size). Prompt: ``In terms of [mass/height/size], it is known that a typical [concept] is [heavy/tall/large]''. Results with $^\dagger$ required manual post-processing of predictions. \label{tabPhysicalProperties}}
\end{table}

\begin{table*}
\centering
\footnotesize
\setlength\tabcolsep{5pt}
\begin{tabular}{lccccccc}
\toprule
& \textbf{ada} & \textbf{babbage} & \textbf{curie} & \textbf{davinci} & \textbf{DeBERTa} &  \textbf{Bi-enc\textsubscript{MSCG+GKB}} & \textbf{Bi-enc\textsubscript{ChatGPT}} \\
\midrule
$p_{\textit{heavy}}$ & 46.6 & \textbf{81.7} & \textbf{64.7} & \textbf{76.8} & \textbf{67.7} & \textbf{69.2} & 42.9\\
$1-p_{\textit{light}}$ & 49.8 & 51.3 & 41.5 & 53.4 & 43.4 &	61.9 &	\textbf{53.4}\\
$p_{\textit{heavy}} \cdot (1-p_{\textit{light}})$ & 46.6 & \textbf{81.7} & \textbf{64.7} & \textbf{76.8} & 47.8 &	\textbf{69.2} &	48.1\\
$p_{\textit{heavy}}/p_{\textit{light}}$ & \textbf{61.4} & 68.3 & 52.1 & 65.9 & 65.2 & 75.7 &46.5\\
\bottomrule
\end{tabular}
\caption{Analysis of alternative strategies for predicting pairwise judgements about mass (accuracy). \label{tabAdditionalMass}}
\end{table*}

\paragraph{Taste Domain}
Table \ref{tabTasteRank} summarises the main results on the taste dataset. For this experiment, each model was used to produce a ranking of the 590 food items, which was then compared with the ground truth in terms of Spearman's rank correlation ($\rho$). For the LMs, we experimented with two prompts. Prompt 1 is of the form ``[food item] tastes [sweet]'', e.g.\ ``apple tastes sweet''. Prompt 2 is of the form ``it is known that [food item] tastes [sweet]''. For the DeBERTa model and the bi-encoders, we verbalised the sweetness property as ``tastes sweet'', and similar for the others. 
The results in Table \ref{tabTasteRank} show a strong correlation, for the GPT-3 models, between model size and performance, with the best results achieved by \emph{davinci}. Similar as was observed for the colour domain by \citet{liu-etal-2022-ever}, there seems to be a qualitative change in performance between LLMs such as davinci and smaller models. While there are some differences between the two prompts, similar patterns are observed for both choices. 
ChatGPT and GPT-4 were not able to provide a ranking of the 590 items, and are thus not considered for this experiment. We also tried ChatGPT on a subset of 50 items, but could not achieve results which were consistently better than random shuffling.

Surprisingly, we find that the DeBERTa model outperforms \emph{davinci} in most cases, and often by a substantial margin. This is despite the fact that the model is 2 to 3 orders of magnitude smaller. Furthermore, we can see a large performance gap between the two variants of the bi-encoder model, with the model trained on ChatGPT examples outperforming \emph{curie}, and even \emph{davinci} in two cases.

Table \ref{tabQualitativeSweetness} presents some examples of the predictions that were made by \emph{davinci} for sweetness (using prompt 1), comparing the ranks according to the ground truth (\emph{gold}) with the ranks according to the \emph{davinci} predictions. The table focuses on some of the most egregious mistakes. As can be seen, \emph{davinci} fails to identify the sweetness of common foods such as chocolate, fruit cake and jam. Conversely, the model significantly overestimates the sweetness of different cheeses and vegetables. 

\paragraph{Physical Properties}
Table \ref{tabPhysicalProperties} summarises the results for the physical properties. For mass, we consider both the problem of ranking all objects, evaluated using Spearman $\rho\%$, and the problem of evaluating  pairwise judgments, evaluated using accuracy. Height and size can only be evaluated in terms of pairwise judgments. To obtain conditional probabilities from the GPT-3 models, we used a prompt of the form ``In terms of [mass/height/size], it is known that a typical [concept] is [heavy/tall/large]''. We also tried a few variants, which performed worse. To compute perplexity scores, for evaluating pairwise judgements, we used a prompt of the form ``[concept 1] is heavier/taller/larger than [concept 2]''. For the baselines, we obtained scores for the properties \emph{heavy}, \emph{tall} and \emph{large}.

The correlation between model size and performance is far from obvious here, except that \emph{ada} clearly underperforms the three larger models. However, among the GPT-3 models, \emph{babbage} actually achieves the best results in several cases. The results based on conditional probabilities are consistently better than those based on perplexity. ChatGPT and GPT-4 were difficult to use with the ranking prompt, as some items were missing, some were duplicated, and many items were paraphrased in the ranking. The results in Table \ref{tabPhysicalProperties} were obtained after manually correcting these issues. With this caveat in mind, it is nonetheless clear that GPT-4 performs exceptionally well in this experiment. In accordance with our findings in the taste domain, DeBERTa performs very well on the height and size properties, outperforming all GPT-3 models by a clear margin. For mass, however, DeBERTa failed completely, even achieving a negative correlation. The bi-encoder models perform well on height and size, although generally underperforming the largest GPT-3 models. For mass, the bi-encoder trained on ChatGPT examples performs poorly, while the model trained on Microsoft Concept Graph and GenericsKB was more robust. It is notable that the results in Table \ref{tabPhysicalProperties} are considerably higher than those obtained by \citet{DBLP:journals/corr/abs-2305-14057} using OPT \cite{DBLP:journals/corr/abs-2205-01068}. For mass, for instance, even the largest OPT model (175B) was not able to do better than random guessing. 

In Table \ref{tabPhysicalProperties}, the pairwise judgments about mass were assessed by predicting the probability of the word \emph{heavy} (for the GPT-3 models) or by predicting the probability that the property \emph{heavy} was satisfied (for the baselines). Another possibility is to use the word/property \emph{light} instead, or to combine the two probabilities. Let us write $p_{\textit{heavy}}$ to denote the probability obtained for \emph{heavy} (i.e.\ the conditional probability of the word, as predicted by the language model, or the probability that the property is satisfied, as predicted by a baseline model), and similar for $p_{\textit{light}}$. Then we can also predict the relative  mass of items based on the value $p_{\textit{heavy}}\cdot (1-p_{\textit{light}})$ or based on the value $p_{\textit{heavy}}/ p_{\textit{light}}$. These different possibilities are evaluated in Table \ref{tabAdditionalMass}. As can be seen, there is no variant that consistently outperforms the others. 

\begin{table}
\footnotesize
\centering
\setlength\tabcolsep{4pt}
\begin{tabular}{l ccccc}
\toprule
& \textbf{Bitter} & \textbf{Sour} & \textbf{Mass} & \textbf{Height} & \textbf{Size} \\
& $\rho$ & $\rho$ & $\rho$ & \textit{Acc} & \textit{Acc}\\
\midrule
Full training & 24.7 & 43.9 & -8.9 & 86.6 & 93.9 \\
Filtered training & 24.8 & 35.0 & 30.7 & 82.0 & 90.8 \\
\bottomrule
\end{tabular}
\caption{Comparison of the DeBERTa model in two settings: the full training setting, where the McRae and CSLB datasets are used for fine-tuning, and a filtered setting, where relevant properties are omitted. \label{tabAnalysisDeBERTa}}
\end{table}

\paragraph{Analysis of Training Data Overlap}
For the baselines, we may wonder to what extent their  knowledge comes from the pre-trained language model, and to what extent it has been injected during the fine-tuning step. For this analysis, we focus in particular on the DeBERTa model, which was fine-tuned on the McRae and CSLB datasets. These datasets indeed cover a number of physical properties, as well as some properties from the taste domain. Table \ref{tabAnalysisDeBERTa} summarises how the performance of the DeBERTa model is affected when removing the most relevant properties from the McRae and CSLB training sets, which we refer to as \emph{filtered training} in the table. For instance, for the property \emph{bitter}, in the filtered setting we omit all training examples involving the properties ``bitter" and ``can be bitter in taste''; for \emph{sour} we remove the properties ``sour'' and ``can be sour in taste''; for \emph{mass} we remove the properties  ``heavy", ``light", ``light weight" and ``can be lightweight";
for \emph{height} we remove the properties ``short", ``can be short", ``tall" and ``can be tall"; and for \emph{size} we remove the properties ``large" and ``small". Note that the McRae and CSLB datasets do not cover any properties that are related to sweetness, saltiness, umami and fattiness. The results in Table \ref{tabAnalysisDeBERTa} show that filtering the training data indeed has an effect on results, although the performance of the model overall remains strong. Interestingly, in the case of \emph{mass}, the filtered setting leads to clearly improved results.

\section{Conclusions}
We proposed the use of a dataset from the taste domain for evaluating the ability of LLMs to learn perceptually grounded representations. We found that LLMs can indeed make meaningful predictions about taste, but also showed that a fine-tuned DeBERTa model, and in some cases even a fine-tuned BERT-large bi-encoder, can outperform GPT-3. The performance of these smaller models crucially depends on the quality of the available training data. For this reason, we explored the idea of collecting training data from ChatGPT, using a new prompting strategy. We complemented our experiments in the taste domain with an evaluation of physical properties, where we achieved considerably better results than those reported in the literature \cite{DBLP:journals/corr/abs-2305-14057}.
%
%
Whereas previous work was essentially aimed at understanding the limitations of language models, our focus was more practical, asking the question: \emph{can high-quality conceptual space representations be distilled from LLMs}? Our experiments suggest that the answer is essentially positive, but that new approaches may be needed to optimally take advantage of the knowledge that can be extracted from such models.


\section*{Limitations}
It is difficult to draw definitive conclusions about the extent to which cognitively meaningful representations can be obtained by querying LLMs. Among others, previous work has found that performance may dramatically differ depending on the prompt which is used; see e.g.\
\cite{liu-etal-2022-ever}. We have attempted to make reasonable choices when deciding on the considered prompts, through initial experiments with a few variations, but clearly this is not a guarantee that our prompts are close to being optimal. However, this also reinforces the conclusion that LLMs are difficult to use directly for learning conceptual spaces. While we believe that taste represents an interesting and under-explored domain, it remains to be verified to what extent LLMs are able to capture perceptual features in other domains. 

\section*{Acknowledgments}
This work was supported by EPSRC grant EP/V025961/1.


\bibliography{anthology,custom}
\bibliographystyle{acl_natbib}

\appendix
\section{Collecting Concept-Property Pairs using ChatGPT} \label{secChatGPTPairs}
To obtain training data for the bi-encoder model using ChatGPT, we used the following prompt:

\begin{quote}
\emph{I am interested in knowing which properties are satisfied by different concepts. I am specifically interested in properties, such as being green, being round, being located in the kitchen or being used for preparing food, rather than in hypernyms. For instance, some examples of what I'm looking for are: \textbf{1. Sunflower, daffodil, banana are yellow 2. Guitar, banjo, mandolin are played by strumming or plucking strings 3. Pillow, blanket, comforter are soft and provide comfort  4. Car, scooter, train have wheels   5. Tree, log, paper are made of wood 6. Study, bathroom, kitchen are located in house}. Please provide me with a list of 50 such examples.}
\end{quote}

We repeated this request with the same prompt around 10 times. After this, we changed the examples that are given (shown in bold above). This process was repeated until we had 287K concept-property pairs. After removing duplicates, a total of 109K such pairs remained. We found it was necessary to regularly change the examples provided to ensure the examples were sufficiently diverse and to avoid having too many duplicates. These examples were constructed manually, to ensure their accuracy and diversity. Asking for more than 50 examples with one prompt became sub-optimal, as the model tends to focus on a narrow set of similar properties when the list becomes too long.

To verify the quality of the generated dataset, we manually inspected 500 of the generated concept-property pairs. In this sample, we identified 6 errors, which suggests that this dataset is of sufficient quality for our intended purpose. Compared to resources such as ConceptNet, the main limitation of the ChatGPT generated dataset is that it appears to be less diverse, in terms of the concepts and properties which are covered. We leave a detailed analysis of this dataset for future work.

\section{Issues with ChatGPT and GPT-4}
For the experiments in Table \ref{tabPhysicalProperties}, we used ChatGPT and GPT-4 to rank 49, 25 and 26 unique objects according to their mass, height and size respectively. The prompt used was as follows: ``Rank the following objects based on their typical [mass/height/size], from [heaviest to the lightest/ tallest to the shortest/ largest to the smallest]", followed by a list of the items. We could not directly evaluate the responses of ChatGPT and GPT-4 because of the following issues:
\begin{itemize}
\item Missing objects: For instance, GPT-4 ranked 48 out of 49 objects and ChatGPT ranked 46 out of 49 objects respectively, according to their mass. 
\item Paraphrasing: While ranking, both GPT-4 and ChatGPT changed some of the the names of the objects. For instance, ``wooden train track" was renamed as "wooden train track piece", ``gardening shears" was renamed as "garden shears".
\item Duplicates: GPT-4 and ChatGPT both occasionally introduced duplicates in the list.
\end{itemize}
To address these issues, we removed the duplicates and appended the missing items at the end of the ranking. We manually corrected the modified names.

\end{document}